HoneyBooBoo:

solve for x 1/4(4x+15)=24

AngeI:

Hello, welcome to QuestionCove o: Do you know how to start the process to solve for x in this scenario?

cpl106:

did you need a decimal or fraction ?

AngeI:

It's the same answer either way but they may convert as needed. I was going to explain it if they didn't reply in a few more minutes so that people who see this post in the future may be helped by it c:

cpl106:

i could give em the answer but i was going to try to explain but i don't think they are paying attention

AngeI:

People often ask questions and leave, I'm unsure why. I'll explain though no worries

AngeI:

First you would have to multiply 1/4 by what is inside the parenthesis 1/4(4x+15) 1/4*4x and 1/4*15 1x+3.75=24 Then you would have to subtract 3.75 from both sides x+3.75-3.75=24-3.75 x=20.25